OpenStudy (anonymous):

-8=-(x+4) solve for x show all work, help anyone ?

OpenStudy (anonymous):

Your goal is to isolate X on one side... In this case divide each side by Negative 1 you are left with 8 = x + 4 now tell me what you would do next?

OpenStudy (anonymous):

//Expand the bracket \[-8=-x-4 \] //Multiply by -1 \[8=x+4\] //Subtract 4 \[8-4 = x\] \[x=4\]